# Tony Lindeberg1    Zahra Babaiee2 and Peyman M. Kiasari2

### v1

@@ 4.3 Implications when using idealized models of “master key filters” in depthwise-separable deep networks
As previously discussed in Section 4.1, it could be interesting to investigate if the addition of an additional filter, corresponding to a discrete approximation of the mixed second-order derivative $\partial_{xy}$ at the joint scale of Filters 1–4, could be advantageous to depthwise-separable network, by making it theoretically possible also to a certain degree of approximation span the space of second-order directional derivative operators according to (130), by in this way enabling formal capture of a significantly richer set of spatial image structures.
As can be seen from the results in Section 3 of estimating the scale values of idealized filter models to approximate the 8 “master key filters” by discrete scale-space operations, the numerical values of the resulting scale estimates can, however, vary notable depending on what criteria are used for computing the scale estimates. For Filter 8, the variability depending on the different criteria is marginal. For Filters 1–6, the variability is, however, significant, because of notable deviations between the idealized receptive field models and the learned filters.
@@
A natural next step to consider would also be to use the proposed structural forms of idealized receptive fields according to the presented methods as initial seeds, to then optimize the choice of parameter settings for the idealized filters in a complementary training stage, in a corresponding way as learning of the scale levels (Pintea et al. (2021, Penaud-Polge et al. (2022, Saldanha et al. 2021, Yang et al. 2023, Basting et al. 2024, Perzanowski and Lindeberg (2025) can be applied to regular Gaussian derivative networks.
@@ 5 Experimental results
To validate the predictive properties of our different idealized filter modelling approaches, we conducted a series of experiments on the ImageNet dataset. Our experimental design follows a similar methodology as described in Babaiee et al. (2025b), where we systematically replace the learned depthwise filters in DS-CNN architectures with idealized approximations derived from our theoretical analysis.
@@ 5.1 Comparative evaluation of idealized filter modelling methods
Table 15 presents the ImageNet Top-1 accuracy results for the ConvNeXt v2 Tiny model when its depthwise filters are replaced with idealized approximations from each of our six methods, without any subsequent fine-tuning of the rest of the parameters.
@@
The results demonstrate that Method B, based on matching discrete weighted variance-based spatial spread measures between idealized receptive field models and learned filters, achieves the highest predictive accuracy at 65.70%. This finding is particularly significant, as it suggests that the discrete modelling approach, which accounts for the inherent discretization effects in the learned filters, provides superior approximation quality compared to methods based on continuous Gaussian derivative models or direct norm minimization.
@@ 5.2 Validation through frozen filter training
Based on the comparative results, we selected Method B for further investigation of its practical applicability in depthwise-separable networks. We conducted experiments where the ConvNeXt v2 Tiny model was initialized with our 8 idealized filters (determined using Method B parameter estimates from Table 8) and trained from scratch on ImageNet with the depthwise filters remaining frozen throughout the 300-epoch training process.
For determining the filter type assignment at initialization, we analyzed the fully trained ConvNeXt v2 Tiny model and classified each depthwise filter according to its closest match among our 8 idealized filter types using the linear approximation procedure described before. These filter type assignments were then fixed throughout the subsequent training process. All the other network parameters were initialized from the FCMAE (fully convolutional masked autoencoder) pre-trained ConvNeXt v2 Tiny weights, following the same initialization procedure as described in the original ConvNeXt v2 paper. For comparative purposes, we also trained a standard ConvNeXt v2 Tiny model using the FCMAE pretrained initialization, following the exact hyperparameters and training procedures from the original code repository. This baseline model achieved 82.7% accuracy and serves as our reference for evaluating the performance impact of our idealized filter constraints.
@@
Original ConvNeXt v2 Tiny 82.7
Frozen 8 master key filters 82.7
Frozen 8 filters from Method B 82.5
With learning of the scale parameters 82.6
Table 16: ImageNet Top-1 accuracy comparison for ConvNeXt v2 Tiny with different filter initialization strategies.
The results reveal a remarkable consistency across different initialization approaches. The model initialized with our Method B idealized filters achieves 82.5% accuracy, representing only a 0.2% decrease compared to the original architecture. This minimal performance degradation, despite using only 8 distinct filter types across all the depthwise layers, provides strong empirical support for our theoretical analysis and demonstrates the effectiveness of our scale-space-based modelling approach.
@@ 5.3 Learning of the scale parameters
To investigate whether the theoretical constraints imposed by our idealized models limit the representational capacity of the network, we conducted an additional experiment where the 8 filter shapes from Method B were frozen, but their scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ were made trainable. This approach maintains the theoretical structure of our scale-space operators, while allowing adaptation to the specific requirements of the layers and architecture.
The model with learnable scale parameters achieved 82.6% accuracy, representing a slight improvement of 0.1% over the fully frozen configuration. This result suggests that while our theoretical parameter estimates provide reasonable initial approximations, gradient-based optimization can yield marginal performance gains.
To understand how the network adapts the scale parameters during training, we analyzed the learned scale parameter distributions across all depthwise layers. Table 17 presents aggregated statistics for the scale parameters $\sigma_{x}$ and $\sigma_{y}$ for each of the 8 filter types after training completion.
@@
Notably, despite these variations in the learned scale parameters—particularly the deviation from our theoretical estimates and the high variance in Filter 8—the overall network accuracy remained essentially unchanged (82.6% vs 82.5%). This suggests that while the specific scale parameters may vary, the fundamental computational structure captured by our 8 idealized filter types is sufficient to maintain performance, indicating that the filter shapes themselves are more critical than their precise parameterization.
@@ 6 Summary and discussion
We have presented an in-depth analysis of different ways of modelling the 8 “master key filters” extracted from the ConvNeXt deep learning architecture by Babaiee et al. (2025a). For this purpose, we have first computed quantitative measures of the spatial extent of the learned filters in terms of both unweighted and weighted spatial spread measures, giving support for modelling the receptive fields of these filters with separable filtering operations along the Cartesian coordinate directions, and using non-centered difference operators for modelling Filters 1–4 and centered difference operators for modelling Filters 5–6, in combination with a spatial smoothing operations with possibly different values of the scale parameters along the different coordinate directions, and possibly different values of the scale parameters for the individual filters.
The presented modelling and analysis results show that the predictive properties of the scale-space framework to derive canonical models for spatial receptive fields in terms of Gaussian smoothing and Gaussian derivatives do qualitatively well generalize from the previously axiomatic formulations regarding the first layer of visual processing, to the qualitative shapes of the receptive fields learned in certain families of depthwise-separable deep networks. Conceptually, these results are also compatible with previous applications of the use Gaussian derivative operations as the computational primitives in deep networks, as developed by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025).
@@
As we have found from the resulting experiments, for some of the learned filters, the scale estimates may vary notable depending on the choice of matching criterion, while for Filter 8 the difference is minor. The main reason for this variability in discrepancy is because of a varying degree of agreement between the idealized models and the learned filters. For judging which of these methods has the best predictive properties with regard to the performance, when used for replacing the computational primitives in depthwise-separable deep networks, this performance should therefore be evaluated experimentally. Specifically, to be less dependent on the choice of mathematical criterion for model fitting, we have proposed to use the presented theory for formulating parameterized primitives for the idealized receptive field models, and then learning these parameters by deep learning, in a corresponding way as learning of the scale levels is used in Gaussian derivative networks.
@@
The experimental results of replacing the 8 “master key filters” in the ConvNeXt v2 Tiny deep network by idealized discrete scale-space filters obtained in this way show that almost as good results are obtained by using the idealized discrete scale-space filters as when using either the 8 “master key filters” or the filters learned by back propagation in the original deep network architecture. Thereby, the results demonstrate that the filters learned in a reduced version of the modern and state-of-the-art deep network architecture ConvNeXt appear to be very similar to the filters obtained by axiomatically determined scale-space theory. Furthermore, we have shown that the additional increase in performance obtained by learning the scale parameters is marginal for the ConvNeXt v2 Tiny model, thus demonstrating the very good predictive ability of Method B when applied to the filters learned in this deep learning architecture.
@@
Notably the amount of spatial smoothing is very small in the idealized models of the 8 “master key filters”. In this modelling work, we have throughout modelled this amount of smoothing with spatial filters based on the discrete analogue of the Gaussian kernel (Lindeberg 1990) combined with small-support difference operators, motivated by the theoretical results in Lindeberg (2024, 2025), that demonstrate that this discretization method is the discretization method that best preserve a transfer of scale-space properties from a continuous to a discrete image domain, and the experimental results in Perzanowski and Lindeberg (2025), that show that this discretization method overall gives the best performance among 5 different discretization methods, when approximating Gaussian derivative operators for application in Gaussian derivative networks.
@@
Concerning the weighted variance-based spread measures for estimating the spatial extents of the kernels, we have in this work throughout used fixed values of the scale parameters for the spatial weighting functions, coarsely determined from the spatial extents of the learned filters as observed by visual inspection. A more general approach in this respect could instead be to vary these scale parameters, or to try to adapt these scale parameters in an iterative manner.
@@
Another conceptual simplification that we have done in the above model fitting schemes is that we have normalized the learned receptive fields that are approximated by first-order derivatives in the horizontal xx–direction to having the same response to the monomial $x$ as a corresponding first-order derivative derivative $g_{x}(x,y;\,\sigma)$. Similarly, we have a pre-processing step to the model fitting stage normalized the learned receptive fields that are approximated by first-order derivatives in the vertical yy–direction to having the same response to the monomial $y$ as a corresponding first-order derivative derivative $g_{y}(x,y;\,\sigma)$. Furthermore, we determined offset constants for Filters 7–8 to minimize their variance-based spatial spread measures, and normalized their responses to a constant unit function to be equal to 1, to make the resulting normalized filters more similar to either a pure spatial smoothing operation or a local sharpening operation. In these ways, we reduced the dimensionality of the following model fitting search, although the resulting model fitting for the $l_{1}$- and $l_{2}$-norm-based modelling schemes may possibly not be as accurate as if it would be extended to also involve a search over the amplitudes of the receptive fields. A possible extension to future work could therefore involve performing such higher-dimensional optimizations over the spaces of the parameters of the idealized receptive fields.
@@
Finally, let us remark that although the work in this paper has been focused on modelling and analyzing the 8 specific “master key filters” extracted by Babaiee et al. (2025a), we propose that a similar methodology could be used for modelling other sets of “master key filters” obtained from other deep networks as well as biological visual receptive fields recorded by neurophysiological measurements.

### v2

@@ 5.3 Implications when using idealized models of “master key filters” in depthwise-separable deep networks
As previously discussed in Section 5.1, it could be interesting to investigate an additional filter, corresponding to a discrete approximation of the mixed second-order derivative $\partial_{xy}$ at the joint scale of Filters 1–4, could be advantageous to depthwise-separable networks. The reason for this is that it would make it theoretically possible also to a certain degree of approximation span the space of second-order directional derivative operators according to (123), and in this way enabling formal capture of a significantly richer set of spatial image structures.
As can be seen from the results in Section 4, of estimating the scale values of idealized filter models to approximate the 8 “master key filters” by discrete scale-space operations, the numerical values of the resulting scale estimates can, however, vary notably depending on what criteria are used for computing the scale estimates. For Filter 8, the variability depending on the different criteria is marginal. For Filters 1–6, the variability is, however, significant, because of notable deviations between the idealized receptive field models and the learned filters.
@@
A natural next step to consider would also be to use the proposed structural forms of idealized receptive fields according to the presented methods as initial seeds, to then optimize the choice of parameter settings for the idealized filters in a complementary training stage, in a corresponding way as learning of the scale levels has been applied in regular Gaussian derivative networks by Pintea et al. (2021), Penaud-Polge et al. (2022), Saldanha et al. (2021), Yang et al. (2023), Basting et al. (2024) and Perzanowski and Lindeberg (2025).
@@ 6 Experimental results
To validate the predictive properties of our different idealized filter modelling approaches, we conducted a series of experiments on the ImageNet dataset. [^9] We used the standard ImageNet-1K dataset, specifically the ILSVRC 2012 (ImageNet Large Scale Visual Recognition Challenge 2012) version, which consists of 1.28 million training images and 50,000 validation images across 1,000 object categories. Our experimental design follows a similar methodology as described in Babaiee et al. (2025b), where we systematically replace the learned depthwise filters in DS-CNN architectures with idealized approximations derived from our theoretical analysis.
@@ 6.1 Comparative evaluation of idealized filter modelling methods
Table 9 presents the ImageNet Top-1 accuracy results for the ConvNeXt V2 Tiny model, when its depthwise filters are replaced with idealized approximations from each of our six methods, without any subsequent fine-tuning of the rest of the parameters.
@@
The results demonstrate that Method B, based on matching discrete weighted variance-based spatial spread measures between idealized receptive field models and learned filters, achieves the highest predictive accuracy at 65.70%. This finding is particularly significant, as it suggests that the discrete modelling approach, which accounts for the inherent discretization effects in the learned filters, provides superior approximation quality, compared to the methods based on continuous Gaussian derivative models or direct norm minimization.
@@ 6.2 Validation through frozen filter training
Based on the comparative results, we selected Method B for further investigation of its practical applicability in depthwise-separable networks. We conducted experiments, where the ConvNeXt V2 Tiny model was initialized with our 8 idealized filters (determined using the estimates of the parameters using Method B in Table 7) and trained from scratch on ImageNet with the depthwise filters remaining frozen throughout the 300-epoch training process.
For determining the filter type assignment at initialization, we analyzed the fully trained ConvNeXt V2 Tiny model and classified each depthwise filter according to its closest match among our 8 idealized filter types using the linear approximation procedure described before. These filter type assignments were then fixed throughout the subsequent training process. All the other network parameters were initialized from the FCMAE (fully convolutional masked autoencoder) pre-trained ConvNeXt V2 Tiny weights, following the same initialization procedure as described in the original ConvNeXt V2 paper. For comparative purposes, we also trained a standard ConvNeXt V2 Tiny model using the FCMAE pretrained initialization, following the exact hyperparameters and training procedures from the original code repository. This baseline model achieved 82.79% accuracy and serves as our reference for evaluating the performance impact of our idealized filter constraints.
@@
Original ConvNeXt V2 Tiny $82.794\pm 0.091$
Frozen 8 master key filters $82.695\pm 0.021$
Frozen 8 filters (Method B) $82.545\pm 0.029$
With learning of the scale parameters $82.609\pm 0.055$
Table 10: ImageNet Top-1 accuracy comparison for ConvNeXt V2 Tiny with different filter initialization strategies. The results are averaged over 3 independent runs, reporting mean $\pm$ standard deviation.
The results reveal a remarkable consistency across different initialization approaches. The model initialized with our Method B idealized filters achieves 82.54% accuracy, representing only a 0.25% decrease compared to the original architecture. This minimal performance degradation, despite using only 8 distinct filter types across all the depthwise layers, provides strong empirical support for our theoretical analysis and demonstrates the effectiveness of our scale-space-based modelling approach.
@@ 6.3 Learning of the scale parameters
To investigate whether the theoretical constraints imposed by our idealized models limit the representational capacity of the network, we conducted an additional experiment, where the 8 filter shapes from Method B were frozen, but their scale parameters $\sigma_{x,i}$ and $\sigma_{y,i}$ were made trainable. This approach maintains the theoretical structure of our scale-space operators, while allowing adaptation to the specific requirements of the layers and the architecture.
Since the modified Bessel functions of integer order, underlying our idealized models of the receptive fields, are not available as full-fledged primitives in PyTorch, as needed for backpropagation of the scale parameters when performing learning of the scale values, we replaced the discrete analogue of the Gaussian kernel in the idealized filter models (95)–(102) with the sampled Gaussian kernel, and combined this kernel with corresponding difference operators according to the methodology in Lindeberg (2025).
The model with learnable scale parameters achieved 82.61% accuracy, representing a slight improvement of 0.06% over the fully frozen configuration. This result suggests that while our theoretical parameter estimates provide reasonable initial approximations, gradient-based optimization can yield marginal performance gains.
To understand how the network adapts the scale parameters during training, we analyzed the learned scale parameter distributions across all the depthwise layers. Table 11 presents aggregated statistics for the scale parameters $\sigma_{x}$ and $\sigma_{y}$ for each of the 8 filter types after training completion.
@@
Notably, despite these variations in the learned scale parameters—particularly the deviation from our theoretical estimates and the high variance in Filter 8—the overall network accuracy remained essentially unchanged (82.61% vs. 82.54%). This suggests that while the specific scale parameters may vary, the fundamental computational structure captured by our 8 idealized filter types is sufficient to maintain good performance, indicating that the filter shapes themselves are more critical than their precise parameterization.
@@ 7 Summary and discussion
We have presented an in-depth analysis of different ways of modelling the 8 “master key filters” extracted from the ConvNeXt deep learning architecture by Babaiee et al. (2025a). For this purpose, we have first computed quantitative measures of the spatial extent of the learned filters in terms of both unweighted and weighted spatial spread measures. This has given support for modelling the receptive fields of these filters with separable filtering operations along the Cartesian coordinate directions, using non-centered difference operators for modelling Filters 1–4 and centered difference operators for modelling Filters 5–6, in combination with spatial smoothing operations. In these spatial smoothing operations, we have allowed for using possibly different values of the scale parameters along the different coordinate directions, and possibly different values of the scale parameters for the individual filters.
The presented modelling and analysis results show that the predictive properties of the scale-space framework, to derive canonical models for spatial receptive fields in terms of Gaussian smoothing and Gaussian derivatives, do qualitatively well generalize from the previously axiomatic formulations regarding the first layer of visual processing, to the qualitative shapes of the receptive fields learned in certain families of depthwise-separable deep networks. Conceptually, these results are also compatible with previous applications of the use Gaussian derivative operations as the computational primitives in deep networks, as developed by Jacobsen et al. (2016), Lindeberg (2021b, 2022), Pintea et al. (2021), Sangalli et al. (2022), Penaud-Polge et al. (2022), Smets et al. (2023), Gavilima-Pilataxi and Ibarra-Fiallo (2023) and Perzanowski and Lindeberg (2025).
@@
As we have found from the resulting experiments, for some of the learned filters, the scale estimates may vary notably depending on the choice of matching criterion, while for Filter 8 the difference is minor. The main reason for this variability in discrepancy is because of varying degrees of agreement between the idealized models and the learned filters. To judge which of these methods should lead to the best predictive properties with regard to the performance, when inserting corresponding idealized filters into depthwise-separable deep networks, this performance should therefore be evaluated experimentally. Specifically, to be less dependent on the choice of mathematical criterion for model fitting, we have proposed to use the presented theory for formulating parameterized primitives for the idealized receptive field models, and then learning these parameters by deep learning, in a corresponding way as learning of the scale levels is used in Gaussian derivative networks.
@@
The experimental results of replacing the 8 “master key filters” in the ConvNeXt V2 Tiny deep network by idealized discrete scale-space filters obtained in this way show that results almost as good are obtained by using the idealized discrete scale-space filters as when using either the 8 “master key filters” or the filters learned by back propagation in the original deep network architecture. Thereby, the results demonstrate that the filters learned in a reduced version of the modern and state-of-the-art deep network architecture ConvNeXt appear to be very similar to the filters obtained by axiomatically determined scale-space theory. Furthermore, we have shown that the additional increase in performance obtained by learning the scale parameters is marginal for the ConvNeXt V2 Tiny model, thus demonstrating the very good predictive ability of Method B when applied to the filters learned in this deep learning architecture.
Notably, the amount of spatial smoothing is very small in the idealized models of the 8 “master key filters”. This use of very fine scale levels is consistent with the results of learning of the scale levels in deep networks based on linear combinations of Gaussian derivative operators, where the results reported in Appendix A.2.1 in the supplementary material of Perzanowski and Lindeberg (2025) also lead to very fine scale levels.
@@
In this modelling work, we have throughout modelled this amount of smoothing with spatial filters based on the discrete analogue of the Gaussian kernel (Lindeberg 1990) combined with small-support difference operators, motivated by the theoretical results in Lindeberg (2024, 2025), that demonstrate that this discretization method is the most suitable one for transferring scale-space properties from a continuous to a discrete image domain, and the experimental results in Perzanowski and Lindeberg (2025), that show that this discretization method overall gives the best performance among 5 different discretization methods, when approximating Gaussian derivative operators for application in Gaussian derivative networks.
@@
Concerning the weighted variance-based spread measures for estimating the spatial extents of the kernels, we used throughout this work fixed values of the scale parameters for the spatial weighting functions, coarsely determined from the spatial extents of the learned filters as observed by visual inspection. A more general approach in this respect could instead be to vary these scale parameters, or to try to adapt these scale parameters in an iterative manner.
@@
Another conceptual simplification, that we have done in the above model fitting schemes, is that we have normalized the learned receptive fields, that are approximated by first-order derivatives in the horizontal xx–direction (or $y$-direction), to having the same response to the monomial $x$ (or $y$) as a corresponding first-order derivative derivative $g_{x}(x,y;\,\sigma)$ (or $g_{y}(x,y;\,\sigma)$). Furthermore, we determined offset constants for Filters 7–8 to minimize their variance-based spatial spread measures, and normalized their responses to a constant unit function to be equal to 1, to make the resulting normalized filters more similar to either a pure spatial smoothing operation or a local sharpening operation.
In these ways, we reduced the dimensionality of the following model fitting search, although the resulting model fitting for the $l_{1}$- and $l_{2}$-norm-based modelling schemes may possibly not be as accurate, as the model fitting could possibly be, if extended to higher-dimensional searches over the model parameters. A possible extension to future work could therefore involve performing such higher-dimensional optimizations over the spaces of the parameters of the idealized receptive fields.
@@
Let us finally remark that although the work in this paper has been focused on modelling and analyzing the 8 specific “master key filters” extracted by Babaiee et al. (2025a), we propose that a similar methodology could be used for modelling other sets of “master key filters” obtained from other deep networks as well as biological visual receptive fields recorded by neurophysiological measurements.